OpenStudy (zmudz):

For what value \(k\) is the following function continuous at \(x=2\)? \(f(x) = \begin{cases} \frac{\sqrt{2x+5}-\sqrt{x+7}}{x-2} & x \neq 2 \\ k & x = 2 \end{cases}\)

ganeshie8 (ganeshie8):

for \(f(x)\) to be continuous at \(x=2\), we must hav \(\lim\limits_{x\to 2}f(x) = f(2)\) that is : \[\lim\limits_{x\to 2}\frac{\sqrt{2x+5}-\sqrt{x+7}}{x-2} = k\] do that conjugate thingy and try evaluating the limit